2b) A bicycle shop sells the following number of bicycles from 1990 to 2000.

Year Number sold

(thousands)

1990 3

1991 3

1992 3

1993 3

1994 6

1995 6

1996 6

1997 6

1998 9

1999 10

2000 12

Compute the first three moving averages of length 3 for the bicycle sales data and

place them in line with the corresponding year.


1
Expert's answer
2021-05-11T14:15:06-0400

The first three moving averages of length 3 for the bicycle sales data  are-


"x_1=\dfrac{3+3+3}{3}=3

\\[9pt]

x_2=\dfrac{3+6+6}{3}=5

\\[9pt]

x_2=\dfrac{6+6+9}{3}=7

\\[9pt]

x_3=\dfrac{10+22}{2}=11"


The table for the first three moving average is-